OpenStudy (anonymous):

June has a set of folding chairs. If she arranges the chairs in 5 rows, she has 2 chairs left over. If she arranges them in 3 rows of the same length, she has 14 left over. How many chairs does she have?

OpenStudy (esshotwired):

Do you have answer choices to choose from?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you help

OpenStudy (esshotwired):

5R+2=S 3R+14=S You would use this equation. Do you know how to solve it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nvm no

OpenStudy (esshotwired):

If you try and solve it, I will tell you if you are right or not. I am not allowed to directly give you the answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it 6

OpenStudy (esshotwired):

No I think you are missing steps or something else.

OpenStudy (esshotwired):

You would change the equations to equal each other so 5R+2=3R+14

OpenStudy (anonymous):

5r+2=3r+14 -2 -2 5r=3r+12 -3r-3r 2r=12 then divide by 2r and I got 6

OpenStudy (esshotwired):

Yes that is right, but 6=R and you have to look for S. So you have to plug in 6 to one of the equations and solve. Then that is your answer.

OpenStudy (anonymous):

where?

OpenStudy (esshotwired):

Well you can plug it into 5R+2=S. 5(6)+2=S 30+2=S 32=S

OpenStudy (esshotwired):

If you do it with the other equation it would still be 32. 3(6)+14=S 18+14=S 32=S

OpenStudy (anonymous):

ohhhh ok that made sense

OpenStudy (esshotwired):

You had the right idea, but just forgot something. Do you understand the steps better now?

OpenStudy (anonymous):

but when you recheck the answer they aren't equal

OpenStudy (anonymous):

on both sides of the equation

OpenStudy (esshotwired):

What do you mean?

OpenStudy (anonymous):

well in my class we are learninghow to sole equations with variables on both sides

OpenStudy (esshotwired):

So what do you mean by they are not equal on both sides?

OpenStudy (anonymous):

nvm I get it

OpenStudy (anonymous):

thx for helping

OpenStudy (esshotwired):

your welcome :)